# Capacity Approximations for Insertion Channels with Small Insertion Probabilities

### v3

@@ V-A Calculation of $H(A^{n},B^{n})$
We define the length of $\bar{Y}(j)$ by $L_{j}$, i.e., $|\bar{Y}(j)|=L_{j}$. We have four cases:
@@
    Case $\mathcal{V}_{1}$: $|\bar{Y}(j)|=|X(j)|$ and $|\bar{Y}(j+1)|=|X(j+1)|$: In this case, we certainly do not have insertion in $X(j)$. Hence, $t_{j}=0$.
@@
    Case $\mathcal{V}_{2}$: $|\bar{Y}(j)|<|X(j)|$: We have a single bit replacement inside $X(j)$. Consider $X(j)$ as a run of 0’s:
@@
    Case $\mathcal{V}_{3}$: $|\bar{Y}(j)|=|X(j)|+1$, $|\bar{Y}(j+1)|=|X(j+1)|$, and there is only a single insertion at the $j$-th and $(j+1)$-th runs (since there will be at most a single insertion in two consecutive runs, we know that there is no insertion at the $(j+2)$-th run).
@@
    $X(j)X(j+1)$ =[0 0 0 1 1 1 1],absentdelimited-[]$0$\ 0\ 1\ 1\ 1\ 1],= [ 0 0 0 1 1 1 1 ] , (139)
    $\bar{Y}(j)\bar{Y}(j+1)$ =[0 0 0 0 1 1 1 1].absentdelimited-[]$0$\ 0\ 0\ 1\ 1\ 1\ 1].= [ 0 0 0 0 1 1 1 1 ] . (140)
@@
    Case $\mathcal{V}_{4}$: $|\bar{Y}(j)|=|X(j)|$, $|\bar{Y}(j+1)|=|X(j+1)|+1$, and there is only a single insertion at the $j$-th and $(j+1)$-th runs.
@@
$H(\check{K}|X^{n},Y)=$ $\frac{\alpha}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)-\epsilon_{2}$ (153)
@@
$0\leq\epsilon_{2}\leq\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right).$ (154)
@@
$H(\check{K}|X^{n},Y)=A_{1}\alpha-\epsilon_{2}+\epsilon_{3},$ (155)
@@
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon},$ (157)
@@
$|H(K|X^{n},Y)-H(\check{K}|X^{n},\check{Y})|\leq 4\alpha^{1+\gamma-\epsilon/2}.$ (158)
@@
$\qquad-\frac{\alpha}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right)\Bigg{)}$
$\qquad+\zeta,$ (159)
@@
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)\leq\epsilon_{1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$
$-2h(z,v^{1},v^{2})\leq\eta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)+2h(z,v^{1},v^{2}),$
@@ V-E Achievability and Converse
The achievability and converse proofs follow a similar structure to those of the first insertion channel model (see Sections IV-E and IV-F). Note that for the converse proof, when calculating $h(z,v^{1},v^{2})$, we observe from (36) and (118) that the bound from (105) can be applied, as $h(z,v^{1},v^{2})\leq h(z,v)$.
@@ VI Conclusions
In this paper, we focused on the capacity of insertion channels with small insertion probabilities, a topic that has been largely overlooked in the literature despite its relevance in practical applications such as DNA coding and storage. Our approach builds upon the series expansion of a previously known capacity point, specifically a channel with no insertions (i.e., zero insertion probability), where the capacity is 1 and is achieved with an i.i.d. Bernoulli$(1/2)$ input distribution. By leveraging this known point, we expect that for small insertion probabilities, the achievable rate will not deviate significantly from this value. With this intuition, we derived the first two terms of the capacity expansion by decomposing the expressions necessary to achieve the maximum rate. Using this methodology, we obtained capacity approximations for two fundamental insertion channel models: 1) the insertion channel with random bit insertions, and 2) the Gallager insertion channel, where a bit is replaced by two random bits. Our results demonstrate that these approximations are achievable by a uniform i.i.d. input distribution up to the first two orders, as expected. For future work, possible extensions include applying similar methods to different channel models or symbol alphabets. For instance, the binary input case could be extended to nonbinary inputs, e.g., 4-ary for DNA storage systems, with deletion and/or insertion channels. Further research could also consider different channel models, such as the absorption channel or the insertion, deletion, and substitution (IDS) channel. Thus, our approach lays the foundation for various extensions and can facilitate capacity analysis for these different models.

### v4

@@ V-A Calculation of $H(A^{n},B^{n})$
Similar to Lemma 8, we proceed from left to right. Although insertions may introduce new runs under our definition of $\bar{Y}$, the indices of $\check{Y}(j)$ and $\bar{Y}(j^{\prime})$ may not coincide; however, the corresponding runs still match and represent the same run segment.
We have four cases:
@@
    Case $\mathcal{V}_{1}$: $|\bar{Y}(j^{\prime})|=|X(j)|$ and $|\bar{Y}(j^{\prime}+1)|=|X(j+1)|$: In this case, we certainly do not have insertion in $X(j)$. Hence, $t_{j}=0$.
@@
    Case $\mathcal{V}_{2}$: $|\bar{Y}(j^{\prime})|<|X(j)|$: We have a single bit replacement inside $X(j)$. Consider $X(j)$ as a run of 0’s:
@@
    Case $\mathcal{V}_{3}$: $|\bar{Y}(j^{\prime})|=|X(j)|+1$, $|\bar{Y}(j^{\prime}+1)|=|X(j+1)|$, and there is only a single insertion at the $j$-th and $(j+1)$-th runs (since there will be at most a single insertion in two consecutive runs, we know that there is no insertion at the $(j+2)$-th run).
@@
    $X(j)X(j+1)$ =[0 0 $1 1],\displaystyle=[0\$0\ 0\ 1\ 1\ 1\ 1], (145)
    $\bar{Y}(j^{\prime})\bar{Y}(j^{\prime}+1)$ =[0 0 0 $1 1].\displaystyle=[0\$0\ 0\ 0\ 1\ 1\ 1\ 1]. (146)
@@
    Case $\mathcal{V}_{4}$: $|\bar{Y}(j^{\prime})|=|X(j)|$, $|\bar{Y}(j^{\prime}+1)|=|X(j+1)|+1$, and there is only a single insertion at the $j$-th and $(j+1)$-th runs.
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\check{K}|X^{n},Y)=$ $\frac{\alpha}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)-\epsilon_{2}$ (159)
@@
$0\leq\epsilon_{2}\leq\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right).$ (160)
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\check{K}|X^{n},Y)=A_{1}\alpha-\epsilon_{2}+\epsilon_{3},$ (161)
@@
$\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon},$ (163)
@@
$\lim_{n\rightarrow\infty}\frac{1}{n}|H(K|X^{n},Y)-H(\check{K}|X^{n},\check{Y})|\leq 4\alpha^{1+\gamma-\epsilon/2}.$ (164)
@@
$\hskip 18.49988pt-\frac{\alpha}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right)\Bigg)$
$\hskip 18.49988pt+\zeta,$ (165)
@@
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)\leq\epsilon_{1}$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$
$-2h(z,v^{1},v^{2})\leq\eta\leq\frac{1}{4}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{4}\alpha^{2}(\mathbb{E}[L_{0}]-1)+2h(z,v^{1},v^{2}),$
@@ V-E Achievability and Converse
The achievability and converse proofs follow a similar structure to those of the first insertion channel model (see Sections IV-E and IV-F).
Lemma 20 (Achievability).
Let $\mathbb{X}^{*}$ be the iid Bernoulli(1/2) process. For any $\epsilon>0$ with the Gallager insertion channel, we have
$C_{2}(\alpha)=1+\alpha\log(\alpha)+G_{2}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}).$ (167)
Proof.
The proof mirrors Section IV-E. Since $\mathbb{X}$ is i.i.d. Bernoulli(1/2)(1/2) of length $n$, and the corresponding output $Y(X^{*,n})$ is also i.i.d. Bernoulli with run length $L_{Y}=n+\text{Bin}(n,\alpha)$, one can repeat the same steps as in Lemma 10 and show that
$H(Y(X^{*,n}))$ $=n(1+\alpha)+\mathcal{O}(\log n).$ (168)
Using the estimate of $H(Y|X^{*,n})$ from Corollary 2, with $P(z,v^{1},v^{2})=\mathcal{O}(\alpha^{2})$ and $\mathbb{E}[L_{0}\log L_{0}]<\infty$, completes the proof. ∎
For the converse, similar to Lemma 11, we start by showing that there is no significant loss if we restrict ourselves to finite-length runs with a maximum length $L^{*}$.
Lemma 21.
Proof.
The proof follows a similar structure to that of [kanoria2010deletion, Lemma III.2] and Lemma 11. Since the steps are provided in detail in Lemma 11, we omit them here and instead provide a sketch of the proof.
We construct $\mathbb{X}_{L^{*}}$ by flipping a bit each time it is the $(\mathcal{S}_{L^{*}}+1)$-th consecutive bit with the same value, with the corresponding output denoted by $Y_{L^{*}}=Y(X^{n}_{L^{*}})$.
We further define $F=F(\mathbb{X},\mathbb{A})$ as a binary vector of the same length as the channel output. The elements of $F$ are set to 1 wherever the corresponding bit in $Y_{L^{*}}$ is flipped relative to $Y$, and 0 otherwise. The entropy $H(F)$ can be computed as:
$H(F)=h(\beta)(1+\alpha)n+\log(n+1),$ (170)
where $L_{F}$ is the length of $F$ and $\beta=\frac{P(L_{0}>L^{*})}{L^{*}}$. The remainder of the proof follows the same steps as in [kanoria2010deletion, Lemma III.2]. ∎
Lemma 22 (Converse).
For any $\epsilon>0$ there exists there exists $\alpha_{0}=\alpha_{0}(\epsilon)>0$ such that the following happens. For any $L^{*}\in\mathbb{N}$ and any $\mathbb{X}\in\mathcal{S}_{L^{*}}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$. If $\alpha<\alpha_{0}(\epsilon)$, then
$I(\mathbb{X})\leq 1+\alpha\log\alpha+G_{2}\alpha+\alpha^{3/2-\epsilon}.$ (171)
Proof.
With a similar approach to that of [kanoria2010deletion, Lemma III.3] and our proof of Lemma 12, we can directly write
$H(Y)\leq n(1+\alpha)+\log(n+1).$ (172)
The proof involves using the lower bound in Corollary 2 together with the function $h(z,v^{1},v^{2})$. Note that $h(z,v^{1},v^{2})\leq h(z)+h(v^{1})+h(v^{2})$. Owing to the similarity between the definitions of $\mathbb{Z}$, $\mathbb{V}^{1}$, and $\mathbb{V}^{2}$ and those of $\mathbb{Z}$ and $\mathbb{V}$ in the simple insertion channel, one can readily show that $h(z,v^{1},v^{2})\leq 3h(z)$. Hence, we have
$h({z},v_{1},v_{2})\leq 3\alpha^{1+\gamma-\epsilon/2}.$ (173)
Also noting that
$\left|\mathbb{E}[\log L_{0}]-\sum_{l=1}^{\infty}2^{-l-1}l\log l\right|=o(\alpha^{1/2-\epsilon}\log L^{*}),$ (174)
the proof follows. ∎
@@ VI Conclusions
In this paper, we focused on the capacity of insertion channels with small insertion probabilities, a topic that has been largely overlooked in the literature despite its relevance in practical applications such as DNA coding and storage. Our approach builds upon the series expansion of a previously known capacity point, specifically a channel with no insertions (i.e., zero insertion probability), where the capacity is 1 and is achieved with an i.i.d. Bernoulli(1/2)(1/2) input distribution. By leveraging this known point, we expect that for small insertion probabilities, the achievable rate will not deviate significantly from this value. With this intuition, we derived the first two terms of the capacity expansion by decomposing the expressions necessary to achieve the maximum rate. Using this methodology, we obtained capacity approximations for two fundamental insertion channel models: 1) the insertion channel with random bit insertions, and 2) the Gallager insertion channel, where a bit is replaced by two random bits. Our results demonstrate that these approximations are achievable by a uniform i.i.d. input distribution up to the first two orders, as expected. For future work, possible extensions include applying similar methods to different channel models or symbol alphabets. For instance, the binary input case could be extended to nonbinary inputs, e.g., 4-ary for DNA storage systems, with deletion and/or insertion channels. Further research could also consider different channel models, such as the absorption channel or the insertion, deletion, and substitution (IDS) channel. Thus, our approach lays the foundation for various extensions and can facilitate capacity analysis for these different models.